OpenStudy (anonymous):

t^2+15t+50 Factor the trinomial. I am not sure how to do this please help.

OpenStudy (anonymous):

what are two numbers a and b so that: a*b=50 a+b=15 ?

OpenStudy (anonymous):

I got 5t^2(3t+10) and that was incorrect

OpenStudy (anonymous):

no that would multiply to 15t^3+50t^2

OpenStudy (anonymous):

so two numbers that multiply to 50 and add to 15 are 5 and 10. So we can factor the trinomial: (x+5)(x+10)

OpenStudy (anonymous):

does this make sense or should I explain why I did this?

OpenStudy (anonymous):

well that answer was not correct

OpenStudy (anonymous):

(t+5)(t+10) is correct

OpenStudy (anonymous):

ok I was putting x sry

OpenStudy (anonymous):

my fault i wrote x too

OpenStudy (anonymous):

t^2+15t+50 You must find a pair of factors of 50 that sum to 15 5 and 10 Now replace the 15t with 5t + 10t t^2+5t + 10t +50 now pair off and factorise... (t^2+5t) + (10t +50) giving ... t(t + 5) + 10(t + 5) answer = (t + 5)(t + 10)